I have tried to solve this problem but when I check it I'm wrong! I'm stumped and don't know what to do here is my work the name of my lesson is solving linear equations using multiplication first:
-2x-5y= 9 -6x - 15y = 27
3x+11y= 4 -6x - 22y = 8
+_____________
7y = 19
y = here is where I got stumped what am I doing wrong???

Found 2 solutions by ReadingBoosters, stanbon:
Answer by ReadingBoosters(3246)

The equations are
-2x-5y=9
3x+11y=4
...
Multiplying the first by 3 and the second by 2
3(-2x-5y=9) ==> -6x - 15y = 27
2(3x+11y=4) ==> 6x + 22y = 8
...
Add the equations
-6x + 6x - 15y + 22y = 27 + 8
7y = 35
y = 5
...
Your error was you have -6 + -6 = 0, when it would equl -12.
The second equation should be multiplied by 2, not -2.
...
11(-2x-5y=9) ==> -22x - 55y = 99
5(3x+11y=4) ==> 15x + 55y = 20
Add the equation
-22x + 15x - 55y + 55y = 99 + 20
-7x = 119
x = -17
...
(-17, 5) is the solution
…………………………….
Have more questions? Get them all answered at once.
HomeworkHelpers@ReadingBoosters.com
Individualized Coaching: www.MyHomeworkAnswers.com
--Reading Boosters

Answer by stanbon(75887)

-2x-5y= 9
3x+11y= 4
-----------
Multiply thru top equation by 3
Multiply thru bottom equation by -2
------------------------------
-6x - 15y = 27
-6x - 22y = -8
----------------------
Subtract to solve for "y":
7y = 35
----
y = 5
---------
Solve for "x" using -2x - 5y = 9
-2x -25 = 9
-2x = 34
x = -17
==================
Cheers,
Stan H.
=========================